OpenStudy (anonymous):

What's is the y value for –3y – 12 = –4

OpenStudy (anonymous):

y= -(8/3)

OpenStudy (cresciez):

Actually y simply equals 6... -3(-6) = - 18 - 12 = -18 + 12 = -4

OpenStudy (cresciez):

Oops sorry remove the negative when you multiply -3 and 6

OpenStudy (cresciez):

But 6 is definitely the answer :P

OpenStudy (anonymous):

well check your answer @Cresciez 6(-3) = 18 18-12 = 6 which would not b the answer not being rude just explaining

OpenStudy (cresciez):

Aw no worries I guess I did a calculation error >->

OpenStudy (anonymous):

-18-12 would be 30 but anyway my first reply was correct. cx